OpenStudy (anonymous):

which of the following is greatest? 6/25 0.12 3/100 9/50

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

woohoo, thanks.

OpenStudy (eujc21):

\[a) 6/25 = 24/100 b) 12/100 c) 3/100 d) 9/50=18/100\]

OpenStudy (anonymous):

thanks!